General Meeting Summary 11/23/16

From cua_phy
Jump to navigation Jump to search

DATA ANALYSES

  • Extraction of pion form factor from kaon experiment E12-09-011 (Salina)
  • Salina showed that E12-09-011 has significant number of pions in acceptance - she's investigating if a precision measurement of the pion form factor would be possible
  • Her plan is to investigate if the kinematic coverage in the azimuthal angle is sufficient with the scheduled configurations - if not, she is planning to estimate the time needed for additional configurations and make a proposal to the spokespeople at the next monthly meeting


DETECTOR UPDATES

  • Aerogel Cherenkov (Arthur)
  • Discussion about options for adding another layer of aerogel tiles - currently have only seven installed
  • main issue: need a different cover for transportation and storage - clearance needed is 3-5 mm
  • options discussed: 1) new custom cover, 2) modification of existing plexiglass and aluminum covers
  • Action item: discuss with Walter what's needed for option (1) and time scale
  • Readout status for cosmic data taking in Hall C (Arthur)
  • Aerogel HV hardware is ready - connected to crate on 2nd floor counting house - network access available
  • Aerogel signal readout can be done through VMEX crate in detector hut and using pre-shower software
  • Trigger for measurements: use a couple of scintillators - 2 channels as done for other trays
  • General status of SHMS DAQ and trigger: hodoscope work still ongoing, full DAQ system will be ready at earliest January 2017
  • NPS PMT shielding studies - scheduling magnetic field exposure for week of 28th
  • Crystal readout (Andres)
  • CUA APD-based readout has been setup - still need to construct shielding, develop a suitable trigger concept, and in long-term think about pre-amp options
  • Greg has arranged for loan of APD-based readout used by HPS to CUA - system comes without trigger
  • Action item: Andres will continue with the setup he built at CUA and discuss with HPS group at JLab about optimizations
  • Crystal measurements (Salina)
  • Richard completed light yield measurements of SICCAS #19 and CRYTUR #01 - data analysis ongoing
  • Salina plans to measure light yield for the same crystals - comparison of these and earlier measurements are expected to shed light on observed discrepancies between light yields measured
  • GluEx detectors (Greg)
  • Update on ongoing tests of cookies (connection between sensor and expansion volume for DIRC) at JLab
  • tests were done with cookies of thicknesses between 1-3 mm and produced by different places - quality still needs to be improved for final application
  • Plans to make transmittance measurements of a number of 2mm thick cookies at CUA in week of 28th - coordinate with Salina


OTHER

  • Discussion about Rachel/Aaron's shipments to CUA


NEXT MEETING: WED. 30 NOV AT 10:00 AM (EST)